# Brenda Prallon22footnotemark: 2    Chen Qiu22footnotemark: 2    Yiwei Sun22footnotemark: 2

### v1

@@ 3 Main Result
Lemma 1 implies that the minimax-regret value of the site selection problem in Definition 1 can be upper bounded by the solution of the uncapacitated $k$-facility location problem. [^6] Uncapacitated means that there is no capacity constraint on the number of clients that each facility can accommodate. See (Williamson and Shmoys, 2011, Chapter 4.5) and Zhang (2007). Just as in the $k$-median problem, there is a set of facilities $\mathcal{F}$ and a set of clients (or cities) $\mathcal{C}$. Now we assume that there is an opening cost $c_{i}\in\mathbb{R}_{+}$ associated with each facility $i\in\mathcal{F}$. The connection cost between facilities and clients is as before. The goal is to open at most $k$ facilities and connect each client to one facility so that total cost is minimized. Thus, the problem
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\left(\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}\right)+\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)\right)$ (15)
can be viewed as a $k$-facility location problem. Just as before, the set of facilities is $\mathcal{S}_{E}$ and the set of clients is $\mathcal{S}_{P}$. The connection cost between sites $s\in\mathcal{S}_{E}$ and $s^{\prime}\in\mathcal{S}_{p}$ is $(C/(2\#\mathcal{S}_{P}))\|X_{s}-X_{s^{\prime}}\|$. The opening cost for any facility $s\in\mathcal{S}_{E}$ is $(B/\#\mathcal{S}_{P})\sigma_{s}$. Minimizing the upper bound in Lemma 1 is thus equivalent to solving the $k$-facility location problem. As shown in Appendix A.1, the upper bound arises by bounding the worst-case sum of regrets across sites by the corresponding sum of worst-case regrets. For selected sites that are both facilities and clients, the worst-case regret is obtained by solving a point-identified treatment choice problem. For policy-relevant sites where no experiment was conducted, it is obtained by solving the partially-identified treatment choice problem in Ishihara and Kitagawa (2021). When $\mathcal{S}_{E}$ and $\mathcal{S}_{P}$ are disjoint, the first component of the upper bound vanishes, and the bound becomes proportional to the solution of the $k$-median problem. ∎
@@
$\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\;\left(\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right).$ (16)
@@
Lemma 2 implies that the minimax-regret value of the site selection problem in Definition 1 can be lower bounded by the solution of the k-median problem:
@@
While the computation of an exact minimax-regret sampling scheme (whether purposive or randomized) appears intractable, our ability to provide nontrivial upper bounds on worst-case regret does not hinge on $C$ being large enough. Indeed, one could always compute such a bound by (i) forcing the treatment decision $T_{s}$ to only depend on site $N_{\mathscr{S}}(s)$ and (ii) computing worst-case expected regret separately across sites, ignoring that the regret-maximizing parameter configurations may be mutually inconsistent across sites. Both manipulations increase regret and therefore define an upper bound. This bound is easy to compute for any given sampling scheme because the induced minimax-regret treatment choice problem was solved in Stoye (2012). For $C$ large enough, it will coincide with the bound in Theorem 1. The caveat is that this bound is nonlinear; hence, while it is easy to compute once, it does not generate a tight connection to the $k$-median problem and will be hard to optimize for large instances of the site selection problem. ∎
@@ 4 Integer Programming and the $k$-median Problem
It is well known (Williamson and Shmoys, 2011, Chapter 7.7, p. 185) that the $k$-median problem in (17) can be formulated as the following linear integer program:
$\min_{\{y_{i},x_{i,j}\}_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}}$ $\sum_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}x_{i,j}\cdot c(j,i)$ (18)
such that $\sum_{i\in\mathcal{S}_{E}}x_{i,j}=1,\;\;\;\;\;\forall j\in\mathcal{S}_{P},$ (19)
$\sum_{i\in\mathcal{S}_{E}}y_{i}\leq k,$ (20)
$0\leq x_{i,j}\leq y_{i},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P},$ (21)
$y_{i}\in\{0,1\},x_{i,j}\in\{0,1\},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}.$ (22)
@@
In this formulation, the choice variables are the $\{0,1\}$-valued variables $y_{i}$ and $x_{i,j}$, for $i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}$. The variable $y_{i}$ indicates whether site $i$ is chosen for experimentation, and $x_{i,j}$ indicates whether experimental site $i$ is the nearest neighbor of policy site $j$. The total number of sites chosen for experimentation is restricted to be no greater than $k$, and site $i$ can only be the nearest neighbor of policy site $j$ if site $i$ is chosen for experimentation. The connection cost between a facility $i$ and a client $j$—denoted $c(j,i)$—is given by $\|X_{i}-X_{j}\|$. [^7] The connection cost in the objective function of the integer program differs from the connection cost in the $k$-median problem in (17) by a constant factor $C/(2\#\mathcal{S}_{P})$, which does not affect the solution of the $k$-median problem. Solving the linear integer program described above is equivalent to solving (17). It is worth highlighting that the program above is linear in the choice variables $x_{i,j}$ and $y_{i}$.
@@
A major advantage of the integer programming formulation of the $k$-median problem is that most ecosystems for scientific computing offer different solvers for linear and nonlinear integer programs. For the applications in this paper, we use the MIP solver in Gurobi (Gurobi Optimization, LLC, 2023) through the Python package gurobipy. The Gurobi software is highly optimized, especially for large-scale problems, and it offers an academic license. Even though the scale of the applications presented in Section 5 is not large enough for the efficiency advantages of Gurobi to become salient over other solvers, we wanted to showcase its ease of use. It also integrates seamlessly with Google Colab, providing us a way to build self-contained and reproducible examples. [^8] https://colab.research.google.com/
@@
The MIP solver uses a version of the branch-and-bound algorithm; see Bertsimas and Weismantel (2005), Chapter 11.1, for a general description of this algorithm. Broadly speaking, this algorithm works by first finding the solution to the linear programming (LP) relaxation of the original integer problem. This is known as the relaxation step. Then, the problem is split into two sub-problems (branching), according to the integer constraints on the solution. This method then gives bounds on the integer solution, and the algorithm is applied recursively until the lower bound and the upper bound converge up to a tolerance parameter. The recursion creates nodes, and there are some strategies to determine which nodes should be explored further; for example, nodes that have integer solutions do not require any more branching.
@@
Gurobi implements several additional steps that help the branch-and-bound algorithm be more efficient. [^9] See https://www.gurobi.com/resources/mixed-integer-programming-mip-a-primer-on-the-basics/. The presolve step reduces the number of effective constraints of the problem by checking if the integer requirement can eliminate some of them. As the name suggests, it is performed before the start of the branch-and-bound algorithm. Cutting planes tightens the feasible region by adding linear inequalities to eliminate fractional solutions; it is performed during the branch-and-bound process, and for this reason the algorithm used by Gurobi can be referred to as a version of a “branch-and-cut” algorithm. Finally, the MIP solver implements several heuristics, for example by rounding the component of a solution that is closest to an integer, fixing it, and hoping the other components will converge to integers quicker.
In principle, one can always solve the $k$-median problem in (17) by enumerating all possible size $k$ subsets of $\mathcal{S}_{E}$. Such an algorithm runs in time proportional to
(#${{\#\mathcal{S}_{E}}\choose{k}}\cdot\#\mathcal{S}_{P}\cdot k\cdot d$
@@
time and therefore scales poorly when the term “#$\#\mathcal{S}_{E}$ choose $k$” is large. However, we are able to evaluate the performance of our preferred solver by brute-force solving smaller but not trivial instances of the problem, notably the entire application in Section 5.2.
@@
Figure 9 in Appendix B.1 shows an example of the output obtained after using the MIP solver in Gurobi to solve the linear integer program for the application in Section 5.2. In this example, the scale of the problem is given by #$\#\mathcal{S}_{E}=\#\mathcal{S}_{P}=15,k=6$, and $d=8$. We defer the details of the application to Section 5.2.
@@ 5.1 Mobile Financial Services in Bangladesh
The experiment was conducted in the Gaibandha district, one of Bangladesh’s poorest regions. It focused on households that had migrant workers in the Dhaka district, the administrative unit in which the capital of Bangladesh is located. Lee et al. (2021) measure several outcomes of both receiving households and sender migrants; see their Figures 3 and 4. To give a concrete example of the measured outcomes, one question of interest is whether families that adopt the mobile banking technology are more (or less) likely to declare that the monga—the seasonal period of hunger in September through November—was not a problem for their household. Table 9, Column 7, p. 61 in Lee et al. (2021) presents results for this specific variable showing that households that used a bKash account in the treatment group are 9.2 percentage points more likely to declare that monga was not a problem.
@@
In Lee et al.’s 2021 words, “[t]he particular nature of our sample potentially limits the external validity” of the analysis. In short, migration corridors differ in characteristics ranging from the distance between the origin and the destination to the cost of living and the average wages. Figure 2 presents a box plot of $d=13$ (standardized) characteristics that Gechter et al. (2024) collected for each of the $41$ migration corridors. We take these corridors to be our potential experimental and policy-relevant sites. That is, $\mathcal{S}_{E}=\mathcal{S}_{P}$, and card($\mathcal{S}_{E}$) = card($\mathcal{S}_{P}$) = 41. Below we present results for $k\in\{1,2\}$.
@@
Figure 2(a) presents a visual comparison of the covariates of Dhaka-Pabna (blue circles) and Dhaka-Gaibandha (pink diamonds), the original site selected by Lee et al. (2021). The covariate values for Dhaka-Gaibandha are slightly outside the interquartile range for at least three covariates: migrant density, mean remittance, and mean household size. In comparison, all but one covariate value for Dhaka-Pabna are within the interquartile range. The figure also shows that two key covariates of Dhaka-Gaibandha are right at the edges of the interquartile range: distance between sites in the corridor (3rd quartile) and mean household income (1st quartile). One might conjecture that the effects of adopting a mobile banking technology to transfer money particularly depend on distance and household incomes, suggesting that the Dhaka-Gaibandha corridor may not be the most representative. [^11] Gechter et al. (2024) suggest that these qualities could explain the large treatment effects found by Lee et al. (2021). Interestingly, Dhaka-Pabna has opposite features: the distance between destination and origin in this corridor is short (1st quartile) and households are relatively better off in terms of income (3rd quartile). The use of the minimax criterion might explain why a corridor with these characteristics may be a good choice for extrapolating experimental results.
Figure 2(b) also presents the covariates of Dhaka-Noakhali (yellow triangles), the migration corridor selected by the Bayesian approach of Gechter et al. (2024). [^12] By construction, the solution of Gechter et al. (2024) depends on the choice of prior. The results herein reported are based on their preferred prior specification; see Section 5.3 p.p.23 in Gechter et al. (2024). We note that for 10 out of the 13 variables that control treatment effect heterogeneity, Dhaka-Noakhali has covariates that are typically outside the interquartile range.
Selected Sites when $k=2$: When $k=2$, the $k$-median solution is to pick Dhaka-Pabna and Dhaka-Pirojpur. The former also solved the $k$-median problem for $k=1$.
Figure 3(a) presents the covariates of Dhaka-Pabna (filled, blue circle) and Dhaka-Pirojpur (hollow, blue circle). While four covariate values of Dhaka-Pirojpur (standard deviation of male migrant wage, standard deviation of household income, price index, and migrant density) are outside the interquartile range, they still appear more central than Dhaka-Gaibandha and Dhaka-Noakhali. However, relative to Dhaka-Pabna, the solution for $k=2$ adds a considerably less central site; Figure 4 illustrates that this site is a good nearest neighbor for some sites that would not otherwise be well matched.
@@
Figure 4 presents a simple visualization of optimal connections between sites under different optimality criteria. Note that the underlying scatter plots in both panels are the same; they visualize the location of corridors in (distance, household income)-space. However, Panel 4(a) indicates which sites were selected through solving the $k$-median problem and which of these sites any other site was matched to, where blue circles represent matches with Dhaka-Pabna and orange crosses represent matches with Dhaka-Pirojpur. Figure 4(a) also presents the sites selected by Gechter et al. (2024), i.e. Dhaka-Magura and Dhaka-Noakhali, both marked with dark blue triangles. They appear close to the sites selected by the $k$-median problem, but in terms of that problem’s criterion function, they only rank $455$ among $820$ candidate solutions and miss the problem’s optimal value by $17\%$. Panel 4(b) similarly visualizes the sites selected by Egami and Lee’s (2024) synthetic purposive sampling approach, i.e. Dhaka-Gopalganj (orange square) and Dhaka-Barguna (blue square). [^14] We generated this figure by using Egami and Lee’s (2024) code on Gechter et al.’s (2024) data. We also visualize how policy sites are matched with experimental sites: Orange crosses correspond to migration corridors that assign all of their weight to Dhaka-Gopalganj; blue circles correspond to migration corridors that assign all of their weight to Dhaka-Barguna; all other sites (gray crossed circles) assign strictly positive weights to both donor sites. This illustrates that these sampling schemes meaningfully differ. This would even be true if the synthetic purposive sampling approach were implemented but forcing degenerate (single donor) matches, because Egami and Lee’s (2024) approach would then reduce to a $k$-mean problem, i.e. using squared Euclidean distance as connection cost.
@@
[…]

### v2

@@ 3 Main Result
Lemma 1 implies that the MMR value of the problem in Definition 1 can be upper bounded by the solution of the uncapacitated $k$-facility location problem. [^7] Uncapacitated means that there is no capacity constraint on the number of clients that each facility can accommodate. See (Williamson and Shmoys, 2011, Chapter 4.5) and Zhang (2007). Just as in the $k$-median problem, there is a set of facilities $\mathcal{F}$ and a set of clients (or cities) $\mathcal{C}$. Now we assume that there is an opening cost $c_{i}\in\mathbb{R}_{+}$ associated with each facility $i\in\mathcal{F}$. The connection cost between facilities and clients is as before. The goal is to open at most $k$ facilities and connect each client to one facility so that total cost is minimized. Thus, the problem
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\frac{B}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}\cap\mathcal{S}_{P}}\sigma_{s}~+~\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)$
can be viewed as a $k$-facility location problem. Just as before, the set of facilities is $\mathcal{S}_{E}$ and the set of clients is $\mathcal{S}_{P}$. The connection cost between sites $s\in\mathcal{S}_{E}$ and $s^{\prime}\in\mathcal{S}_{p}$ is $(C/(2\#\mathcal{S}_{P}))\|X_{s}-X_{s^{\prime}}\|$. The opening cost for any facility $s\in\mathcal{S}_{E}$ is $(B/\#\mathcal{S}_{P})\sigma_{s}$. Minimizing the upper bound in Lemma 1 is thus equivalent to solving the $k$-facility location problem. As shown in Appendix A.1, the upper bound arises by bounding the worst-case sum of regrets across sites by the corresponding sum of worst-case regrets. For selected sites that are both facilities and clients, the worst-case regret is obtained by solving a point-identified treatment choice problem. In this case, the optimal treatment rule for each $s\in\mathcal{S}_{P}\cap\mathscr{S}$ is to simply treat the whole population if $\hat{\tau}_{s}\geq 0$. For policy-relevant sites where no experiment was conducted, it is obtained by solving the partially-identified treatment choice problem in Ishihara and Kitagawa (2021), for which the optimal treatment rule for each site $s\in\mathcal{S}_{P}\backslash\mathscr{S}$, when $C>C^{*}$, depends only on the estimate from its nearest neighbor and may randomize (Yata, 2021; Montiel Olea et al., 2025). See Appendix A.1 for the exact form of these treatment rules. When $\mathcal{S}_{E}$ and $\mathcal{S}_{P}$ are disjoint, the first component of the upper bound vanishes, and the bound becomes proportional to the solution of the $k$-median problem. ∎
@@
$\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\geq\;\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|.$
@@
Lemma 2 implies that the MMR value of the problem in Definition 1 can be lower bounded by the solution of the k-median problem
@@
Our results extend to non-equal weights on policy relevant sites with minor modifications. Denote by $\omega_{s}>0$ a known weight for each site $s\in\mathcal{S}_{P}$. Without loss of generality, assume that $\sum_{s\in\mathcal{S}_{P}}\omega_{s}=\#\mathcal{S}_{P}$. We may define weighted welfare as
$\mathcal{W}^{\omega}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\omega_{s}\tau(X_{s})\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})],$
with weighted regret
$\mathcal{R}^{\omega}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\omega_{s}\tau(X_{s})\left(\mathbf{1}\{\tau(X_{s})\geq 0\}-\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right).$ (11)
Then, the minimax-regret optimal purposive sampling scheme and treatment rule are as in Definition 1 but with the modified $\mathcal{R}^{\omega}$. Inspecting proofs of Lemmas 1 and 2, we find that the modified problem can be approximated by the alternative $k$-median problem
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\sum_{s\in\mathcal{S}_{P}}\omega_{s}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|,$
i.e., the connection cost between client $j$ and facility $i$ is now $\omega_{j}\|X_{j}-X_{i}\|$. Let $\omega_{P}:=\max_{s\in\mathcal{S}_{P}}\left\{\omega_{s}\right\}$. Theorem 1 can then be modified as follows: For the same $C^{*}$ defined there and for any $\mathscr{S}\in\mathcal{A}(k)$,
$\Bigg{|}\left(\inf_{T\in\mathcal{T}_{\mathscr{S}}^{1/2}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}^{\omega}(T,\mathscr{S},\tau)\right)-\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\cdot\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\omega_{s}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\Bigg{|}$
is at most
$B\cdot\omega_{P}\cdot\sigma_{E}\cdot\frac{\min\{\#\left(\mathcal{S}_{E}\cap\mathcal{S}_{P}\right),k\}}{\#\mathcal{S}_{P}}.$
We finish this section by mentioning that—even when experimental and policy-relevant sites are disjoint and Assumptions 1-2 hold—we were not able to find an exact MMR sampling scheme (whether purposive or randomized) for all possible values of $C$ (the Lipschitz constant that controls treatment effect heterogeneity). One challenge is that, absent enough treatment effect heterogeneity, the optimal treatment rule for each policy-relevant site will tend to aggregate information from multiple experimental sites, and will generally trade-off the representativeness of a candidate set of experimental sites (as measured by the distance between covariates) against the precision of the corresponding treatment effect estimators. To see this, consider the extreme case where there is no site-level heterogeneity; i.e., $C=0$. In this scenario, any experimental site provides an unbiased estimator for the true treatment effect in any policy-relevant site. Consequently, making policy choices for the policy-relevant sites based only on the information available for its nearest-neighbor tends to be suboptimal. One can show that for each set of experimental sites, $\mathscr{S}$, and for each policy-relevant site, $s\in\mathcal{S}_{P}$, the optimal treatment rule at the policy-relevant site $s$ is a weighted linear combination of the treatment effects in the experimental sites:
$T_{s}(\widehat{\tau}_{\mathscr{S}})=\mathbf{1}\left\{w_{\mathscr{S}}^{\top}\widehat{\tau}_{\mathscr{S}}\geq 0\right\},\textrm{ where }w_{\mathscr{S}}=\bigl{(}w_{\mathscr{S}_{1}},w_{\mathscr{S}_{2}},\ldots,w_{\mathscr{S}_{\#\mathscr{S}}}\bigr{)}^{\top}\text{ and }w_{\mathscr{S}_{i}}=\frac{1/\sigma_{\mathscr{S}_{i}}^{2}}{\sum_{i=1}^{\#\mathscr{S}}\left(1/\sigma_{\mathscr{S}_{i}}^{2}\right)}.$
That is, the optimal treatment assignment aggregates an inverse-variance weighted average of all experimental sites—which provides an efficient estimator, given the available information, of the treatment effect at any policy-relevant site $s$. Thus, the optimal sampling scheme would simply choose the $k$ sites in $\mathcal{S}_{E}$ that lead to the smallest-variance estimator of the true treatment effect.
In the more general case, where $C>0$, it is possible to characterize the optimal treatment rule at each policy-relevant site; see Montiel Olea et al. (2025). However, this site-by-site solution only provides an upper bound on the problem’s MMR value. Moreover, implementing this upper bound has the additional challenge that the ex-ante variance of the hypothetical estimate from each experimental site needs to be known.
@@
In contrast, a strength of our $k$-median proposal is that it does not require ex-ante knowledge of these variances. In addition, for general values of $C$, our approach also provides a nontrivial upper bound on the problem’s MMR value: Just i) force the treatment decision $T_{s}$ to only depend on site $N_{\mathscr{S}}(s)$ and ii) compute worst-case expected regret separately across sites, ignoring that the regret-maximizing parameter configurations may be mutually inconsistent across sites. Both manipulations increase regret and therefore define an upper bound, which is furthermore easy to compute for any given sampling scheme because the induced MMR treatment choice problem was solved in Stoye (2012). For $C$ large enough, the bound will recover Theorem 1. The caveat is that, in general, choosing the sites that minimize this bound will require brute-force enumeration.
@@ 4 Integer Programming and the $k$-median Problem
The $k$-median problem in (10) can be formulated as the following linear integer program (Williamson and Shmoys, 2011, Chapter 7.7, p. 185):
$\min_{\{y_{i},x_{i,j}\}_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}}$ $\sum_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}x_{i,j}\cdot c(j,i)$
such that $\sum_{i\in\mathcal{S}_{E}}x_{i,j}=1,\;\;\;\;\;\forall j\in\mathcal{S}_{P},$
$\sum_{i\in\mathcal{S}_{E}}y_{i}\leq k,$
$0\leq x_{i,j}\leq y_{i},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P},$
$y_{i}\in\{0,1\},x_{i,j}\in\{0,1\},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}.$
@@
Here, the choice variables are the binary-valued $y_{i}$ and $x_{i,j}$ (for $i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}$); $y_{i}$ indicates whether site $i$ is chosen for experimentation, and $x_{i,j}$ indicates whether experimental site $i$ is used for policy choices at the policy-relevant site $j$. The total number of sites chosen for experimentation cannot exceed $k$, and site $i$ can only be used for making policy choices at policy site $j$ if site $i$ is indeed chosen for experimentation. The connection cost between facility $i$ and client $j$ is $c(j,i):=\|X_{i}-X_{j}\|$. [^8] The connection cost in the objective function of the integer program differs from the connection cost in the $k$-median problem in (10) by a constant factor $C/(2\#\mathcal{S}_{P})$, which does not affect the solution of the $k$-median problem. Solving the linear integer program described above is equivalent to solving (10).
@@
A major advantage of this integer programming formulation is that most ecosystems for scientific computing offer different solvers for linear and nonlinear integer programs. For the applications in this paper, we use the MIP solver in Gurobi (Gurobi Optimization, LLC, 2023) through the Python package gurobipy. The Gurobi software is highly optimized, especially for large-scale problems, and it offers an academic license. Even though the scale of the applications presented in Section 5 is not large enough for the efficiency advantages of Gurobi to become salient over other solvers, we wanted to showcase its ease of use. It also integrates seamlessly with Google Colab, providing us a way to build self-contained and reproducible examples. [^9] https://colab.research.google.com/
@@
The MIP solver uses a version of the branch-and-bound algorithm; see Bertsimas and Weismantel (2005), Chapter 11.1, for a general description of this algorithm. Broadly speaking, this algorithm works by first finding the solution to the linear programming (LP) relaxation of the original integer problem. This is known as the relaxation step. Then, the problem is split into two sub-problems (branching) according to the integer constraints. This gives bounds on the integer solution; the algorithm is applied recursively until the lower bound and the upper bound converge up to a tolerance parameter. The recursion creates nodes, and there are some strategies to determine which nodes should be explored further; for example, nodes that have integer solutions do not require any more branching.
@@
Gurobi implements several additional steps that help the branch-and-bound algorithm be more efficient. [^10] See https://www.gurobi.com/resources/mixed-integer-programming-mip-a-primer-on-the-basics/. The presolve step reduces the number of effective constraints of the problem by checking if the integer requirement can eliminate some of them. As the name suggests, it is performed before the start of the branch-and-bound algorithm. Cutting planes tightens the feasible region by adding linear inequalities to eliminate fractional solutions; it is performed during the branch-and-bound process, and for this reason, the algorithm used by Gurobi can be referred to as a version of a “branch-and-cut” algorithm. Finally, the MIP solver implements several heuristics, for example by rounding the component of a solution that is closest to an integer, fixing it, and hoping the other components will converge to integers more quickly.
@@
In principle, one can always solve (10) by enumerating all possible size $k$ subsets of $\mathcal{S}_{E}$. Such an algorithm runs in time proportional to ${{\#\mathcal{S}_{E}}\choose{k}}\cdot\#\mathcal{S}_{P}\cdot k\cdot d$ and therefore scales poorly when ${\#\mathcal{S}_{E}}\choose{k}$} is large. However, we are able to evaluate the performance of our preferred solver by brute-force solving smaller but nontrivial instances of the problem, notably the entire application in Section 5.2.
@@
Figure 9 in Appendix B.1 shows an example of the output obtained after using the MIP solver in Gurobi to solve the linear integer program for the application in Section 5.2. In this example, the scale of the problem is given by $\mathcal{S}_{E}=\#\mathcal{S}_{P}=15,k=6$, and $d=8$. We defer the details of the application to Section 5.2.
@@ 5.1 Mobile Financial Services in Bangladesh
The experiment was conducted in the Gaibandha district, one of Bangladesh’s poorest regions. It focused on households that had migrant workers in the Dhaka district, the administrative unit in which the capital of Bangladesh is located. Lee et al. (2021) measured several outcomes of both receiving households and sender migrants; see their Figures 3 and 4. To give a concrete example of the measured outcomes, one question of interest is whether families that adopt the mobile banking technology are more (or less) likely to declare that the monga, the seasonal period of hunger in September through November, was not a problem for their household. Lee et al. (2021) (Table 9, Column 7) present results for this specific variable showing that households that used a bKash account in the treatment group are 9.2 percentage points more likely to declare that monga was not a problem.
@@
In Lee et al.’s 2021 words, “[t]he particular nature of our sample potentially limits the external validity” of the analysis. In short, migration corridors differ in characteristics ranging from distance between origin and destination to cost of living and average wages. Figure 2 presents a box plot of $d=13$ (standardized) characteristics that Gechter et al. (2024) collected for each of the $41$ migration corridors. We take these corridors to be our potential experimental and policy-relevant sites. That is, $\mathcal{S}_{E}=\mathcal{S}_{P}$, and $\mathcal{S}_{E}=\#\mathcal{S}_{P}=41$. Below we present results for $k\in\{1,2\}$.
@@
Figure 2(a) presents a visual comparison of the covariates of Dhaka-Pabna (blue circles) and Dhaka-Gaibandha (pink diamonds), the original site selected by Lee et al. (2021). The covariate values for Dhaka-Gaibandha are slightly outside the interquartile range for at least three covariates: migrant density, mean remittance, and mean household size. In comparison, all but one covariate value for Dhaka-Pabna are within the interquartile range. The figure also shows that two key covariates of Dhaka-Gaibandha are right at the edges of the interquartile range: distance between sites in the corridor (3rd quartile) and mean household income (1st quartile). One might conjecture that the effects of adopting a mobile banking technology to transfer money particularly depend on distance and household incomes, suggesting that the Dhaka-Gaibandha corridor may not be the most representative. [^12] Gechter et al. (2024) suggest that these qualities could explain the large treatment effects found by Lee et al. (2021). Dhaka-Pabna has opposite features: The distance between destination and origin in this corridor is short (1st quartile) and households are relatively better off in terms of income (3rd quartile). The use of the minimax criterion might explain why a corridor with these characteristics may be a good choice for extrapolating experimental results.
Figure 2(b) also presents the covariates of Dhaka-Noakhali (yellow triangles), the migration corridor selected by the Bayesian approach of Gechter et al. (2024). [^13] By construction, the solution of Gechter et al. (2024) depends on the choice of prior. The results herein reported are based on their preferred prior specification; see Section 5.3 p.p.23 in Gechter et al. (2024). For 10 out of the 13 variables that control treatment effect heterogeneity, Dhaka-Noakhali has covariates that are typically outside the interquartile range.
Selected Sites when $k=2$: When $k=2$, the $k$-median solution is to again pick Dhaka-Pabna and additionally Dhaka-Pirojpur. Figure 3(a) presents the covariates of Dhaka-Pabna (filled, blue circle) and Dhaka-Pirojpur (hollow, blue circle). While four covariate values of Dhaka-Pirojpur are outside the interquartile range, they still appear more central than Dhaka-Gaibandha and Dhaka-Noakhali. However, relative to Dhaka-Pabna, the solution for $k=2$ adds a considerably less central site; Figure 4 illustrates that this site is a good nearest neighbor for some sites that would not otherwise be well matched.
@@
Figure 4 presents a simple visualization of optimal connections between sites under different optimality criteria. The underlying scatter plots in both panels are the same; they visualize the location of corridors in (distance, household income)-space. Panel 4(a) indicates which sites were selected through solving the $k$-median problem and which of these sites any other site was matched to, where blue circles represent matches with Dhaka-Pabna and orange crosses represent matches with Dhaka-Pirojpur. Figure 4(a) also presents the sites selected by Gechter et al. (2024), i.e. Dhaka-Magura and Dhaka-Noakhali, both marked with dark blue triangles. They appear close to the sites selected by the $k$-median problem, but in terms of that problem’s criterion function, they only rank $455$ among $820$ candidate solutions and miss the problem’s optimal value by $17\%$. Panel 4(b) similarly visualizes the sites selected by Egami and Lee’s (2024) synthetic purposive sampling approach, i.e. Dhaka-Gopalganj (orange square) and Dhaka-Barguna (blue square). [^15] We generated this figure by using Egami and Lee’s (2024) code on Gechter et al.’s (2024) data. We also visualize how policy sites are matched with experimental sites: Orange crosses correspond to migration corridors that assign all of their weight to Dhaka-Gopalganj; blue circles correspond to migration corridors that assign all of their weight to Dhaka-Barguna; all other sites (gray crossed circles) assign strictly positive weights to both donor sites. This illustrates that these sampling schemes meaningfully differ. This would even be true if the synthetic purposive sampling approach were implemented but forcing degenerate (single donor) matches, because Egami and Lee’s (2024) approach would then reduce to a $k$-mean problem, i.e. using squared Euclidean distance as connection cost. Note also that we decided not to report how policy sites are matched with experimental sites in the Bayesian solution of Gechter (2024). This is simply because, under standard Gaussian process priors, the posterior mean for the treatment effect at each policy relevant site uses the information available for all experimental sites. It would be interesting to analyze whether treatment assignment rules that only use the information of some experimental sites could arise from sparsity-inducing priors such as those discussed in Datta et al. (2016).
@@
Indeed, Figure 5(a) compares time to solve the $k$-median problem for $k\in\{1,\ldots,10\}$ using a) the integer program formulation of the $k$-median problem in Section 4 and b) brute-force enumeration. The MIP solver in Gurobi solves all instances of the problem to provable optimality in less than one second each. In contrast, brute-force enumeration takes approximately $5$ hours for k=10k=10.161616This was run in a Windows XPS with $10$ cores and 3232GB of RAM using R. The code is parallelized, and to avoid memory issues when $k>8$, it runs a C++ function in the background to evaluate one combination at a time. This ensures that we are giving brute-force enumeration the best chance of success.